1. A property owner owes 6.0 mills in school taxes, 12.0 mills in county taxes, and 4.0 mills in municipal taxes and qualifies for a $40,000 homestead exemption. How much is property tax, assuming an assessed value of $160,000?

2. A street is to be paved and gutters installed in front of your property. The city assesses property owners 75% of the cost of such improvements, which is estimated at $60 per running foot. Your property has 100 feet of frontage on one side of the street. How much will be your special assessment?

3. Calculate the tax rate for a city based on the following estimates:

Forecast expenditures         $70,000,000

Non-property tax income        15,000,000

Total property values        3,000,000,000

Tax exempt properties        100,000,000
Answer #1

1. Taxable value of property = $160,000 - $40,000 = $120,000

Value of mills in per 1,000

School taxes = 6/1000 = 0.006

Country taxes = 12/1000 =0.012

Municipal taxes = 4/1000= 0.004

Total Property tax = $120,000 * 0.006 + $120,000 * 0.012 + $120,000 * 0.004 = $2,640.

Please upvote , if found the answer useful.
For any clarification, Feel free to reach at comment Section